# Adaptive Group Policy Optimization: Towards Stable Training and Token-Efficient Reasoning

### v1

@@ Abstract
Since DeepSeek-R1 popularized, Group Relative Policy Optimization (GRPO) has become the core part of Reasoning LLMs training. However, we find some deficiency that influences RL stability and inference efficiency. Thus, we propose Adaptive Group Policy Optimization (AGPO) which contains two simple but effective modifications: a revised advantage estimation method to mitigate zero-variance situations; a length-based reward, incentivizing the model to avoid overthinking. The experiments demonstrate our methods achieve more stable training and comparable or superior performance with significantly fewer tokens in reasoning steps.
This is an unfinished version and will be updated. We aim to share some findings.
@@ 1 Introduction
Large Language Models (LLMs)Bommasani et al. (2021); Wei et al. (2022); Zhao et al. (2023) have achieved impressive performance through extensive pre-training and fine-tuning processes. However, effectively generating desired model responses often necessitates aligning outputs with specific downstream tasks and human preferencesWang et al. (2023); Wolf et al. (2023).
For alignment challenges, reinforcement learning from human feedback (RLHF)Bai et al. (2022); Kaufmann et al. (2023) is introduced as a prominent post-training strategy, adopted by notable LLMs including GPT-4, Claude, Gemini, and DeepSeek. They have explored various optimization techniques such as Proximal Policy Optimization (PPO)Schulman et al. (2017). Recently, to significantly reduce computational and memory overhead associated with PPO, DeepSeek eliminated the value model and proposed Group Relative Policy Optimization (GRPO)Guo et al. (2025), which achieves high computational efficiency and excellent reasoning performance, surpassing other open-source models ranging from 7B to 70B.
@@
(1) Vanished Advantage Signal: GRPO computes the normalized advantage $A_{i}=\frac{r_{i}-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}}$ for each sampled output $i$ where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of group rewards. However, when all rewards within a group are identical, this normalized numerator and denominator approach 0, causing the advantage signal to vanish, which potentially stalling training.
(2) Inefficient CoT Length: While lengthy As GRPO lacks mechanisms to discourage excessively long chain-of-thought (CoT), models tend to produce overly verbose explanations. A refined approach that rewards concise, effective reasoning is essential to improve token efficiency.
@@
  • Training Efficiency: By identifying the limitations of GRPO’s advantage, we introduce a modified advantage estimation method that addresses zero-variance scenarios, ensuring continuous and effective learning.
  • Token Efficiency: We incorporate a self-adaptive length reward into RL training. Compared with GRPO baselines, our approach achieves comparable performance with significantly fewer response tokens.
@@ 2.1 Policy Gradient
Policy gradient method is one of the most fundamental RL algorithm which directly model and optimize the policy. For any differentiable policy, the policy gradient is :
@@
Where advantage $A_{t}$ is the most crucial part for policy gradient method, which defines how much better a specific action $a_{t}$ is compared to average action give a state $s_{t}$.
@@ 2.2 Proximal Policy Optimization (PPO)
PPO (Schulman et al., 2017) is one of policy gradient methods which uses clipped surrogate objective for policy optimization. Specifically, it maximize the following objective:
@@ 3 Adaptive Group Policy Optimization
During RL training with the GRPO algorithm, we find two useful modifications to improve efficiency and performance.
3.1 Modification 1: Corner case of Advantage estimation
It is easy to find that if the rewards of the group are all equal, like all correct or all wrong, then the advantages become all zero. Therefore, the corresponding sample in the batch has no effect in the training. As model training progresses, this situation increases. We believe that it is beneficial to incorporate the group accuracy signal into the training. The new advantage function is as follows:
$A_{i}=\left\{\begin{array}[]{ll}1,&r_{mean}=r_{max}\\ -1,&r_{mean}=r_{min}\\ \frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})},&\text{otherwise}\end{array}\right.$ (6)
By replacing the original GRPO with Equation 6, the objective maintains all information in the batch, including the directions that policy models follow to reason better.
3.2 Modification 2: Length Reward
It is noticeable that reasoning models often have overthinking problem (Kumar et al., 2025). To mitigate it and improve token efficiency, we applied the length reward described in (Team et al., 2025) by adding it to the accuracy reward, but the performance got severe decline during the training. Hence, common length rewards are not suitable for GRPO settings. We design a simple but effective method that dramatically reduces response length while achieving same model effect. The length reward is defined as follows:
$r_{len}(i)=\left\{\begin{array}[]{ll}0,&\text{if }\text{mean}(r_{acc})=r_{max}\\ \lambda,&\text{otherwise}\end{array}\right.$ (7)
where,
$\lambda=\left\{\begin{array}[]{ll}0,&r_{acc}(i)=r_{min}\\ 1-\frac{len(o_{i})-min_{i}len(o_{i})}{max_{i}len(o_{i})-min_{i}len(o_{i})},&r_{acc}(i)=r_{max}\end{array}\right.$
In Equation 7, we set length reward as 0 when outputs are all correct to avoid negative weighted mean advantages. We only use $i\in correct\_set{\{o_{1},o_{2},\cdots,o_{G}\}}$ when calculate $max_{i}len(o_{i})$ and $min_{i}len(o_{i})$ in Equation 8. The final reward is $r=r_{acc}+\gamma r_{len}$.
@@ 4 Experiment
We conduct a few experiments for evaluating how these modifications affect the RL training of Reasoning models.
@@ 4.1 Implementation Details
We use Qwen2.5-Math-7B as the base model. All experiments are conducted on the MATH train set (Hendrycks et al., 2021) and tested on MATH500 (Lightman et al., 2023) with the metric of Pass@1. VeRL (Sheng et al., 2024) is utilized to perform RL training on the base model with a mini batch size of 32 and a learning rate of $61e-$ 6. The number of group rollout is 5. As for reward settings, the $r_{min}$ and $r_{max}$ for accuracy reward are set to 0 and 1 respectively. $\gamma$, the coefficient of length reward, is $0.1$.
4.2 Training Efficiency
Figure 1: Policy loss curves of GRPO and GRPO m1(modification 1)
Figure 2: Test set performance of GRPO and GRPO m1
After applying modification 1, we observe a more consistent and continuously decreased loss compared to the original GRPO from Figure 1. It provides the training more sufficient learning signals from batch data. Figure 2 also shows that the performance of GRPO-m1 is more stable at the early stage and exceeds the performance of GRPO most of the time. They achieve 78.2 and 76.2 at the best checkpoint respectively.
4.3 Token Efficiency
Model Overall Level 1 Level 2 Level 3 Level 4 Level 5
Qwen2.5-Math-7B 61.0 (620) 86.0 (299) 82.2 (454) 67.6 (584) 57.0 (632) 37.3 (853)
Qwen2.5-Math-7B-GRPO 77.2 (640) 97.7 (354) 86.7 (437) 89.5 (521) 76.4 (707) 55.7 (863)
Qwen2.5-Math-7B-AGPO 77.2 (463) 97.7 (231) 88.9 (327) 91.4 (369) 72.4 (507) 55.7 (671)
Table 1: Comparison of performance and token consumption on MATH500. We record results of overall and different difficulty levels with the format of pass@1 (average response tokens).
We add modification 2 on top of GRPO-m1 in order to further improve token efficiency. As Table 1 shows, our method dramatically reduces response length from 640 to 463 while keeping on-par performance. There is a $22\%$-$35\%$ improvement in token efficiency among difficulty levels. For validating fine-grained results of difficulty levels, we use evaluation codes of OpenR (Wang et al., 2024) after training finished, so the accuracy values may have a slight difference than those in subsection 4.2.
@@ 4.3 Token Efficiency
We will experiment on different base models and datasets in future to validate universality of our methods. More ablation studies around modifications will be taken as well. Besides, we plan to generalize modification 1 to all groups in a batch and adjust pre-defined advantage values of corner cases.

### v2

@@ Abstract
Since DeepSeek-R1 popularized, Group Relative Policy Optimization (GRPO) has become the core part of training Reasoning LLMs. However, we find some deficiency that influences RL stability and inference efficiency, like zero-variance in advantage estimation. Thus, we propose Adaptive Group Policy Optimization (AGPO) which contains a simple but effective modification: a revised objective function to mitigate training fluctuation and zero advantage. The experiments demonstrate our method achieves more stable training and superior performance with significantly fewer tokens in reasoning steps.
@@ 1 Introduction
Large Language Models (LLMs)Bommasani et al. (2021); Wei et al. (2022); Zhao et al. (2023) have achieved impressive performance through extensive pre-training and post-training processes. However, effectively generating desired model responses often necessitates aligning outputs with specific downstream tasks and human preferencesWang et al. (2023); Wolf et al. (2023).
For alignment challenges, reinforcement learning from human feedback (RLHF)Bai et al. (2022); Kaufmann et al. (2023) is introduced as a prominent post-training strategy, adopted by notable LLMs including GPT-4, Claude, Gemini, and DeepSeek. They have explored various optimization techniques such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Direct Preference Optimization (DPO) Rafailov et al. (2023). Recently, to significantly reduce computational and memory overhead associated with PPO, DeepSeek eliminated the value model and proposed Group Relative Policy Optimization (GRPO)Guo et al. (2025), which achieved high computational efficiency and excellent reasoning performance, surpassing other open-source models ranging from 7B to 70B.
@@
(1) Confusing Training Signal: Negative losses happen in RL training, but in this scenario it is not always beneficial. Higher group accuracy may have lower advantage thus influencing loss estimation. Besides, when all rewards within a group are identical, the normalized advantage approaches 0, causing the loss signal to vanish, which potentially stalling training.
(2) Inefficient CoT Length: Since GRPO lacks mechanisms to discourage excessively long chain-of-thought (CoT), models tend to produce overly verbose explanations. A refined approach that rewards concise and effective reasoning is essential to improve token efficiency.
@@
  • Training Efficiency: By identifying the limitations of GRPO’s advantage, we introduce an adaptive loss function that addresses negative loss and zero advantage scenarios, ensuring continuous and effective learning.
  • Token Efficiency: Our adaptive loss implicitly improves token efficiency. Compared with GRPO baselines, our approach achieves better performance with significantly fewer response tokens.
@@ 2.1 Policy Gradient
Policy gradient method is one of the most fundamental RL algorithm that directly model and optimize the policy. For any differentiable policy, the policy gradient is :
@@
Where advantage $A_{t}$ is the most crucial part for policy gradient method, which defines how much better a specific action $a_{t}$ is, compared to average action given a state $s_{t}$.
@@ 2.2 Proximal Policy Optimization (PPO)
PPO (Schulman et al., 2017) is one of the policy gradient methods which uses clipped surrogate objective for policy optimization. Specifically, it maximize the following objective:
@@ 3 Adaptive Group Policy Optimization
During RL training with the GRPO algorithm, we find a useful modification to improve performance and efficiency.
3.1 Adaptive Loss
It is easy to find that if the rewards of the group are all equal, like all correct or all wrong, then the advantages become all zero. Therefore, the corresponding sample in the batch has no effect in the training. As model training progresses, this situation increases. We believe that it is helpful to mask out these positions from the loss calculation.
Another issue is entropy collapse occurred in GRPO training. We observed the same phenomenon as in other works, like DAPO (Yu et al., 2025). The entropy drops quickly leading the policy model to give up exploring. However, we consider it as the problem of objective design rather than clip ratio. Current negative loss value tends to be overconfident since a group with higher accuracy may have higher deviation of rewards which may cause lower advantage compared to other groups in batch. We propose an adaptive loss to tackle these issues:
$L_{(q)\sim P(Q)}=\left\{\begin{array}[]{ll}masked,&\text{if }\{o_{i}\}_{i=1}^{G}\text{ all correct or wrong}\\ max(0,-J_{\text{GRPO}}(\theta)),&\text{otherwise}\end{array}\right.$
Usually in RL training, the model is updated through a mini batch size of training data. Therefore, the loss mask in Equation 6 takes place in the mean operation of losses of questions $q$ in the batch. Negative loss values are also clipped for the mean loss of the batch in order to control fluctuation and maintain a suitable training level. We believe the clip avoids the model getting stuck in a local optimum.
By replacing the original objective, our method focuses on useful information in the batch and normalizes the loss for stable training, which would bring performance improvement and token efficiency.
@@ 4 Experiment
We conduct a few experiments for evaluating how our method affects the RL training of reasoning models.
@@ 4.1 Implementation Details
We use Qwen2.5-7B and Qwen2.5-14B as the base models. All experiments are conducted on our curated dataset, which is constructed by mixing data from MATH train set (Hendrycks et al., 2021) and DAPO train set (Yu et al., 2025). For MATH train set, only data where difficulty levels are greater than or equal to 3 are selected. Similarly, for DAPO train set, data are retained only if solution rates achieved by Qwen3-32B model (Yang et al., 2025) are fall between 0.5 and 0.8 inclusively. These filtering techniques are to ensure the difficulty distribution across the obtained dataset is balanced.
@@
VeRL (Sheng et al., 2024) is utilized to perform RL training with a train batch size of 32, a PPO mini batch size of 8 and a learning rate of $61e-$ 6. The number of group rollout is 8. Temperature for generation is set to 1. As for reward settings, the $r_{correct}$ and $r_{wrong}$ for accuracy reward are set to 0 and 1 respectively. Trained checkpoints that achieve best performance on MATH-500 (Lightman et al., 2023) with the metric of Pass@1 are selected for further evaluation with respect to token efficiency.
It is worth noting that KL divergence penalty is not applied for all experiments. This is based on the observation that model distribution can vary significantly compared to reference model during long CoT training. Therefore, removing KL divergence has been adopted as common practice in the domain.
4.2 Main Results
Model MATH-500 (Pass@1)
Qwen2.5-7B 44.0
Qwen2.5-7B-GRPO 73.2
Qwen2.5-7B-AGPO (w/o loss mask) 73.4
Qwen2.5-7B-AGPO (w/o loss clip) 73.0
Qwen2.5-7B-AGPO 74.6
Qwen2.5-14B 59.8
Qwen2.5-14B-GRPO 75.4
Qwen2.5-14B-AGPO (w/o loss mask) 75.0
Qwen2.5-14B-AGPO (w/o loss clip) 75.2
Qwen2.5-14B-AGPO 77.2
Table 1: Performance of different RL techniques on MATH-500.
Model Average Response Tokens
Qwen2.5-7B 571
Qwen2.5-7B-GRPO 699
Qwen2.5-7B-AGPO 533
Qwen2.5-14B 772
Qwen2.5-14B-GRPO 574
Qwen2.5-14B-AGPO 521
Table 2: Average response length of different RL techniques on MATH-500.
Table 1 shows the performance of different models on the benchmark. Both GRPO and AGPO acquire huge performance gains compared with the base models. As for Qwen2.5-7B experiments, we observe a clear improvement on MATH-500 from 73.2 to 74.6 at the best checkpoint. Qwen2.5-14B experiments also induce similar conclusion that the adaptive loss further refines the model by 1.8 percentage. It is obvious that both loss clip and loss mask are important for our method. If we train without the mask, the performance drops even lower than that of GRPO for all model sizes and it is same for loss clip.
Table 2 illustrates the token efficiency of different models on the benchmark. Qwen2.5-7B-AGPO only takes 533 tokens on average for solving MATH-500 problems while Qwen2.5-7B-GRPO consumes 699 tokens that is $31\%$ higher. In terms of 14B models, our AGPO also ranks first in token efficiency which uses 521 tokens averagely.
4.3 Training Dynamics
Figure 1: Actor entropy curves of GRPO and AGPO for Qwen-2.5-7B
Figure 2: Actor entropy curves of GRPO and AGPO for Qwen-2.5-14B
We examine several variations of training metrics after applying AGPO.
As shown in Figure 1 and Figure 2, a significant enhancement in entropy is observed for both Qwen2.5-7B and Qwen2.5-14B actor models after the application of AGPO. This observation can be attributed to the lower loss clip operation in AGPO, where only positive training loss from below-average actions is maintained while negative training loss is overconfident and controlled in the gradient update. Consequently, the probability distribution is drifted upward asymmetrically by adequate gradient without being drifted downward meanwhile, which manifests higher measured entropy during training. The higher entropy eventually facilitates the generation of more diversified samples within the batch, which is essential for large-scale RL training.
Figure 3: Response length curves of GRPO and AGPO for Qwen2.5-7B
Figure 4: Response length curves of GRPO and AGPO for Qwen2.5-14B
Figure 5: Reward score curves of GRPO and AGPO for Qwen2.5-7B
Figure 6: Reward score curves of GRPO and AGPO for Qwen2.5-14B
We also find substantial reductions in response length while the training steps increase, for both models with AGPO as Figure 3 and Figure 4 show. Meanwhile, comparable accuracy performance is maintained on training set as shown in Figure 5 and Figure 6. This phenomenon can be attributed to masking loss operation in AGPO, since the effect of loss mask on response length is clearly shown in Figure 3 and Figure 4. In this operation, the losses of all correct or wrong groups are masked, enhancing accurate estimation of average loss of the batch. The correction consequently amplifies gradient update towards correct direction. Therefore, in combination with the empirical observation, we believe the loss mask implicitly serves as length-based reward to constrain response length.
5 Conclusion
In this work, we propose a novel method, AGPO, to train a more powerful reasoning model. Our adaptive loss, including loss clip and mask, demonstrates noticeable improvement on both model performance and inference efficiency. Also, our method helps to avoid entropy collapse while training. We want to do more designs about the adaptive loss as future directions. For example, the current loss can be normalized to non-negative values by exponential equations.
@@ 5 Conclusion
We will experiment on more kinds of base models and datasets in future to validate universality of our method. More ablation studies around modifications will be taken as well. It is also uncertain if our approach can produce effects together with other tricks proposed by different GRPO refinements.